# Contextual Embedding Learning to Enhance 2D Networks for Volumetric Image Segmentation

### v1

@@ 3.5 Efficiency Comparison
We further compared the baseline 2D networks, corresponding 3D versions, and the CE-enhanced 2D networks in terms of parameter amount and floating point operations (FLOPs). As shown in Table 3, the CE block only increased the number of parameters by about 10M compared with the original 2D networks, which is about half the model size of the corresponding 3D versions. In addition, the CE-enhanced networks are more lightweight compared with the corresponding 3D networks in terms of FLOPs. The comparison results in Table 3 demonstrate the proposed CE block is a memory-efficient solution to enhance the 2D networks for volumetric segmentation.
@@ 4 Discussion
The automatic segmentation of organs in volumetric medical images, such as MRI and CT scans, plays an important role in computer-assisted diagnosis and interventions [45, 46, 43]. Conventional 2D convolutional networks treat volumetric images as a stack of 2D slices thus cannot learn spatial information in the third dimension [44]. The 3D CNNs are beneficial to extract richer volumetric information, but often require high computation cost and severe GPU memory consumption. To alleviate these issues, the proposed contextual embedding strategy can convert conventional 2D segmentation network consisting of the encoder-decoder architecture into a pseudo-3D segmentation network that captures the slice-wise context, which solves the issue in a memory-friendly way that the segmentation results of 2D networks are discontinuous and incomplete in the 3D space. The results on prostate MRI images and liver CT images show that the proposed CE block can effectively enhance the performance of U-net and DeepLabV3 on the volumetric segmentation. In some cases, the CE-enhanced 2D networks even generated more accurate segmentation metrics than the corresponding 3D networks did. From the visual comparison in Fig. 7 and Fig. 8, it is obviously that comparing with the conventional 2D networks, the CE-enhanced methods can provide more reliable results in the coronal and sagittal views. All above results corroborate that the CE block has strong capability to utilize the slice-wise context for the accurate volumetric segmentation.

### v2

@@ 3.5 Efficiency Comparison
We further compared the baseline 2D networks, corresponding 3D versions, and the CE-enhanced 2D networks in terms of parameter amount and floating point operations (FLOPs). As shown in Table 3, it can be observed that compared to the 2D backbone networks, the parameter number of the proposed CE Block is relatively small (about 0.2 M). In addition, the CE-enhanced U-Net and DeepLabV3 are more lightweight compared with the corresponding 3D networks in terms of FLOPs. The comparison results in Table 3 demonstrate the proposed CE block is a memory-efficient solution to enhance the 2D networks for volumetric segmentation.
@@ 4 Discussion
The automatic segmentation of organs in volumetric medical images, such as MRI and CT scans, plays an important role in computer-assisted diagnosis and interventions. For example, it facilitates the calculation of the required radiotherapy dose for specific region-of-interests. In addition, the segmentation results can be further used to assist mask/surface-based medical image registration. Conventional 2D convolutional networks treat volumetric images as a stack of 2D slices thus cannot learn spatial information in the third dimension. The 3D CNNs are beneficial to extract richer volumetric information, but often require high computation cost and severe GPU memory consumption.
To alleviate above issues, the proposed contextual embedding strategy can convert conventional 2D segmentation network consisting of the encoder-decoder architecture into a pseudo-3D segmentation network that captures the slice-wise context, which solves the issue in a memory-friendly way that the segmentation results of 2D networks are discontinuous and incomplete in the 3D space. Our assumption is that in the embedding space, pixels of the same type exhibit similarity, reflected in the similarity of distances calculated by Eq (1). The resulting distance map positively impacts the segmentation results by providing valuable feedback to help the network distinguish between pixels of the same type. Conversely, if the network incorrectly assigns pixels of different types to the same category, it adversely affects the results. These feedbacks contribute to optimizing the network. In summary, dividing pixels into the same category or different categories is a process of network adaptation. As shown in Fig. 9, segmentation of image $S_{n}$ without using CE Block yields the result $P_{n^{\prime}}$, where the red region is the prediction and the white mask denotes ground truth. In contrast, $D_{n}$ is the distance map between $S_{n}$ and $S_{n-l}$, and $P_{n-l}$ is the refined prediction of image $S_{n-l}$. It can be observed that the upper right part of the liver in the distance map suggests similarity to the previous frame in the embedding space, indicating that this region should be segmented as part of the liver. We refine $P_{n^{\prime}}$ using $D_{n}$ to obtain better result $P_{n-CE}$ for $S_{n}$.
Figure 9: An example to visualize the distance map in the embedding space enhancing the segmentation.
@@
While our method demonstrates improved results with the original 2D network, it does not optimize the backbone network itself. Therefore, the segmentation performance may still subject to the backbone architecture. As shown in Fig. 9, although the proposed CE block can refine the segmentation results, there still exists under-segmentation region due to the inherent capability of the backbone network. Additionally, in this study, our method has only been validated on two single-organ segmentation tasks. Its performance across a wider range of anatomical structures, or its adaptability to different imaging modalities (i.e., ultrasound images) could be interesting future directions.
@@ 5 Conclusion
Acknowledgements
This work was supported in part by the National Natural Science Foundation of China under Grant 62071305, in part by the Guangdong-Hong Kong Joint Funding for Technology and Innovation under Grant 2023A0505010021, in part by the Guangdong Basic and Applied Basic Research Foundation under Grant 2022A1515011241, and in part by the Shenzhen Municipal Scheme for Basic Research under Grant JCYJ20210324100208022.